OpenStudy (mendicant_bias):

(ODE) Problem regarding the Existence of a Unique Solution theorem, prompt posted below shortly.

OpenStudy (mendicant_bias):

http://i.imgur.com/dFmZsKy.png

OpenStudy (mendicant_bias):

The first one is something that I literally could copy-paste from the textbook, but I struggle with paraphrasing it in a meaningful way because I don't quite get the way it works.

ganeshie8 (ganeshie8):

this is hard existence and uniqueness sounds like real analysis to me lol

OpenStudy (mendicant_bias):

Lol. I'm trying to find where it talks about it in my book, it's hidden away somewhere.

OpenStudy (mendicant_bias):

http://i.imgur.com/CCOdRUZ.png

OpenStudy (mendicant_bias):

Looking for equation (2) really quick.

OpenStudy (mendicant_bias):

http://i.imgur.com/EdXZjCA.png

OpenStudy (mendicant_bias):

Lol....god, these guys are too freaking lazy to just rewrite the equation, let me go find equation (1) now.

OpenStudy (mendicant_bias):

http://i.imgur.com/E0w3sn9.png

OpenStudy (mendicant_bias):

Alright, so that's just an IVP.

OpenStudy (mendicant_bias):

So it's the uniqueness/existence of a unique solution to an IVP.

OpenStudy (mendicant_bias):

Just gonna try and solve the second part, that should give me a better feel for how the theorem works.

OpenStudy (mendicant_bias):

\[y \frac{dy}{dx}=xy+1\]

OpenStudy (mendicant_bias):

IVP needs to be in the general form of, \[\frac{dy}{dx}=f(x,y);\]Putting this problem into that form: \[\frac{dy}{dx}=\frac{xy+1}{y}\]

ganeshie8 (ganeshie8):

that equation cannot be solved analytically with any of the standard known methods

OpenStudy (mendicant_bias):

Alright, now, uh, something something, rectangular region, partial with respect to y, MAGIC

OpenStudy (mendicant_bias):

y not

OpenStudy (dan815):

the main thing is thatu cannot have 2 different slopes at a point

OpenStudy (dan815):

solution slope curves cannot intersect

OpenStudy (mendicant_bias):

How do I know that this isn't solvable from looking at the problem, or that it does not have a solution via that theorem?

ganeshie8 (ganeshie8):

we need to "just show" that an unique solution exists using the existence and uniqueness thm we don't need to find the actual solution

OpenStudy (mendicant_bias):

Yeah, so trying to show that...

OpenStudy (mendicant_bias):

Needs to have (5,3) in it, and first partial with respect to y must be continuous, along with the function itself, in that region.

ganeshie8 (ganeshie8):

ohk.. im still trying to interpret the statemetn of the thm

OpenStudy (mendicant_bias):

Yeah, same here

OpenStudy (mendicant_bias):

Well, alright, I'm just going to select a 2x2 region for R and see if it is continuous inside that region. Let's see. \[R=[2,4]\times[4,6]\]

OpenStudy (mendicant_bias):

Denominator is nonzero, everything looks sensible, no discontinuities over that given region for either x or y to my understanding, so that part of the theorem I think is fulfilled.

ganeshie8 (ganeshie8):

f(x,y) is continuous everywhere except at y=0

OpenStudy (mendicant_bias):

Yeah, and that's not in our region, so we're good AFAIK

ganeshie8 (ganeshie8):

same with \(\frac{\partial f}{\partial y} = \frac{-1}{y^2}\)

OpenStudy (mendicant_bias):

Need to check \[\frac{\partial f}{\partial y}\]

OpenStudy (mendicant_bias):

Oh, ok

OpenStudy (mendicant_bias):

Alright, lol, welp, hopefully we properly applied the theorem, and it looks like that's that, heh

OpenStudy (mendicant_bias):

Wait a minute

OpenStudy (mendicant_bias):

Got my point wrong, it was (-5,3), not (5,3)

OpenStudy (mendicant_bias):

It still should work, but with a different region R, I believe

ganeshie8 (ganeshie8):

yeah pick the region in second quadrant

OpenStudy (mendicant_bias):

\[R=[-4,-6]\times[2,4]\]

OpenStudy (mendicant_bias):

http://gifsec.com/wp-content/uploads/2014/02/boss-gifs.gif

ganeshie8 (ganeshie8):

doesnt work

OpenStudy (mendicant_bias):

Wot

ganeshie8 (ganeshie8):

you have (0,0) in your region

OpenStudy (mendicant_bias):

Lemme, lemme draw the thing I'm imagining

OpenStudy (mendicant_bias):



OpenStudy (mendicant_bias):

I think we're confused about how I notated the square, the left hand bracket was the x side, from negative four to negative six, and the right hand bracket was the y side/length, from two to four

ganeshie8 (ganeshie8):

Got you !!!

OpenStudy (mendicant_bias):